# Cryptocurrency as an Investable Asset Class: Coming of Age

### v2

@@ 3.4 Fact 4: How to be “smart” in crypto—crypto-size, crypto-momentum and crypto-value
Fact 4: A small number of smart beta strategies—based on crypto-size, crypto-momentum, and crypto-value—span the cross-section of cryptocurrency returns.
@@
A central question in asset pricing is whether systematic variation in expected returns can be explained by a parsimonious set of factors. In equities, substantial academic endeavor—most prominently by Eugene Fama and Kenneth French (e.g., Fama & French (1992))—established this approach, and factor models play a core role in modern finance practice and theory from risk management of large banks to niche hedge fund strategies. For cryptocurrencies, Liu et al. (2022) establish that a three factor model consisting of the aggregate cryptocurrency market, size, and momentum (C-3 factor model), captures a large fraction of the cross-section of expected returns. Liu et al. (2021) show that adding a value factor (C-4 factor model) further strengthens the explanatory power. We revisit this evidence using our updated sample and show that a small number of predictors drive the cross-section of cryptocurrency returns over the longer sample and post-2020.
@@
We construct smart beta strategies by sorting coins into quintile portfolios based on three established significant predictors: market capitalization (CSIZE), past two-week returns (CMOM), and the price-to-new-address ratio (CVALUE). [^3] Studies based on smaller samples of coins have also proposed additional predictors of cryptocurrency returns in the cross-section, including liquidity (e.g., Bianchi et al. 2022), market segmentation (e.g., Borri & Shakhnov 2022), and higher-order moments (e.g., Borri et al. 2022b; Borri & Santucci de Magistris 2022). The return to each predictor is measured as the long-short spread between the highest and lowest quintiles (Portfolio 5 minus Portfolio 1).
@@
Table 2, Panels A and B, report average excess returns to the quintile portfolios and the corresponding long-short strategies for the full sample and the post-2020 subsample. All three predictors generate statistically significant long-short returns in the full sample. The average long-short returns based on size and value are significantly negative and the average long-short return based on momentum is significantly positive, consistent with Liu et al. (2022). Cryptocurrency value is the weekly quintile portfolio of each bracket based on the price-to-new address ratio. High bracket is the low-value coins and low bracket is the high-value coins. The sorting is done in week $t$ and the returns are the value-weighted average in the following week.
@@
The C-4 structure of the crypto-market thus parallels the equity market, where a small number of equity-market factors—most notably size, value, momentum, and profitability—dominate the cross-section of expected returns. For example, Asness et al. (2013) present evidence of value and momentum in equities, bonds, currencies, and commodities. The C-4 factor model of Liu et al. (2022) and Liu et al. (2021) is the cryptocurrency counterpart of the classic 4-factor model of Fama & French (1992) and Carhart (1997) for equities. It is important to note that the predictors in cryptocurrency markets are only partly analogous to those in equity markets. CSIZE and CMOM closely resemble their equity counterparts, although crypto momentum is measured over a much shorter horizon (previous two weeks, versus the standard 12-to-2 months in equities). In contrast, there is no direct analogue to book value (Shiller 1981) in cryptocurrencies, so we follow Liu et al. (2021) and proxy CVALUE with the price-to-new-address ratio. [^4] Liu et al. (2021) show that the number of new addresses captures the network effects of cryptocurrencies and serves as a measure of their fundamental value.
Establishing the small factor structure of the cryptocurrency returns, in our view, is not only important from purely the empirical asset pricing point of view but also since it provides a set of stylized low dimensional features that capture the main insights of several important theoretical models of cryptocurrency (Cong et al. 2021b, Sockin & Xiong 2023a, b).
A related important point is that in addition to determining which betas are smart, it is equally as important to establish which betas from the factor zoo are not “smart”. Specifically, Liu et al. (2022) and Liu et al. (2021) show that—besides crypto-specific size, momentum, and value—most of the smart beta strategies from the canonical factor zoo in the equity market (Feng et al. 2020) either do not generate significant hedged long-short strategy returns or are subsumed by the C-4 factor model. In Table 3, we study long-short strategies based on these signals post–2020. These strategies include portfolios based on long-horizon momentum, volatility, volume, and beta. We find that these long-short strategies continue not to generate significant average returns in the recent sample. [^5] In a recent study, Fieberg et al. (2024) examine 20,736 research designs for 43 sorting variables in the cryptocurrency market. They find that the most prominent cryptocurrency factors, such as size and momentum, remain consistently robust across various specifications, but the others do not. These results give evidence that the cross-section of cryptocurrency returns can be summarized by a small number of factors.
@@
Notes: This table reports the mean quintile portfolio returns and long-short portfolio return for crypto smart beta strategies based on insignificant predictors. The predictors are past 12-week and 24-week return (MOM12 and MOM24), past 24-week return standard deviation (VOL), past 24-week dollar trading volume (VOLUME), coin market beta (CMKT BET) and stock market beta (MKT BETA) estimated using past 24-week returns. The long-short portfolio is long portfolio 5 and short portfolio 1 for all predictors. The mean returns are the time-series averages of weekly value-weighted portfolio excess returns. The post-2020 sample is the sample since Jan 1, 2020. $t$-statistics in parentheses are Newey-West adjusted. *, **, and *** denote significance at the 10%, 5% and 1% levels.
3.5 Fact 5: Mind the Jumps—large, sudden price moves and “common disasters”
Fact 5: Jumps are frequent and account for a large share of variation of cryptocurrency returns, far more than in equities. Thus, disasters are not rare but common.
@@ 3.5 Fact 5: Mind the Jumps—large, sudden price moves and “common disasters”
Leibniz wrote that natura non facit saltus—nature makes no jumps. In economics and finance, this often means that most changes occur gradually, with only occasional discontinuities. Indeed, jumps are rare in many asset classes, and this key insight leads Black & Scholes (1973), in their celebrated option pricing paper, to model prices as Brownian motions. [^6] Merton (1976) is aware of potential jumps in equity prices and include the feature in his paper. Rare disasters are also used as a central ingredient to explain risk-premia (e.g., Rietz (1988); Barro (2006)). Cryptocurrencies are different: jumps, and hence disasters, are not “rare disasters” but are “common disasters.” Specifically, Borri & Santucci de Magistris (2022) show that 12.8% of Bitcoin’s daily variability can be attributed to jumps—large, sudden price moves—and that higher-order moments such as skewness and kurtosis explain the risk premium on Bitcoin. [^7] See early evidence from Scaillet et al. (2020) similarly documenting very frequent jumps in Bitcoin. Fact 5 revisits and extend this finding, documenting the frequency and size of jumps in cryptocurrency and showing that those are far more relevant than for the U.S. stock market.
Table 4 reports the frequency and magnitude of jumps in Bitcoin and Ethereum—the two largest cryptocurrencies by market capitalization—compared with U.S. stocks. Jumps are identified as the difference between the daily realized variance of Barndorff-Nielsen & Shephard (2002) and the bi-power variation of Barndorff-Nielsen & Shephard (2006), a jump-robust measure of integrated variance. The two realized measures are constructed from tick-level data aggregated at the 5-minute frequency. For U.S. equities, the same procedure is applied, and values represent cross-sectional averages across the constituents of the Dow Jones Industrial Average. Only jumps that are statistically significant at the 0.1% confidence level according to the BNS test of Barndorff-Nielsen & Shephard (2006) are retained. For Bitcoin and Ethereum, the probability of a significant jump on a given day is around 40 percent in the pre–2020 sample and 15–20 percent in the post–2020 period. In sharp contrast, the probability of a daily jump in U.S. stocks is below 10 percent before 2020 and falls to about 3 percent thereafter. Thus, jumps are far more frequent in cryptocurrencies than in equities. Conditional on a jump occurring, however, the share of total return variation explained by the jump is broadly similar across asset classes, particularly in the post–2020 period. This shows that what sets cryptocurrencies apart is the high frequency of jumps—contrary to Leibniz’s dictum that natura non facit saltus, cryptocurrencies are markets where saltus sunt regula (jumps are the rule), rather than the exception.
These results show that higher-order moments are central to understanding cryptocurrency risk. Equity and bond markets follow relatively smooth dynamics with only rare discontinuities. Cryptocurrency markets are characterized by frequent and sizable jumps that are a defining feature of their return distribution. This prevalence of jumps, and the corresponding role of higher-order moments, means significant challenges for risk management and portfolio allocation (see, e.g., Borri 2019). The implication is that we need empirical asset-pricing strategies that explicitly accommodate higher-order moments and nonlinearities (see, e.g., Kozak et al. 2020, Chetverikov et al. Forthcoming, Borri et al. 2025).
@@
This table reports the probability of observing a statistically significant jump (P(J)), the number of such events (N(J)), and quantiles of the ratio of jump variation to realized variance (J/RV), conditional on a significant jump, for Bitcoin, Ethereum, and U.S. stocks. Significant jumps are identified using the BNS test with a 0.1% significance level. For U.S. equities, values are cross-sectional averages across the constituents of the Dow Jones Industrial Average. Data are from January 1, 2015 to March 31, 2025. The pre-2020 sample is the sample ending Jan 1, 2020. The post-2020 sample is the sample since Jan 1, 2020.
3.6 Fact 6: Few factors, higher orders—rather than machine learning: why less is more
Fact 6: In cryptocurrency, higher-order terms of the C-4 model accounts for approximately 25% of the cross-section of returns and captures most of the insights of machine learning methods.
A growing literature in finance uses machine learning methods, such as neural networks, to investigate the cross-section of returns (see, for example, Gu et al. 2020). Neural networks have become extremely successful in many prediction problems and moreover have the universal approximation property (Cybenko 1989; Hornik et al. 1989)—a sufficiently large neural network can approximate any continuous function—and thus, in principle, capture any features of the data. These ideas have been present in the machine learning literature for several decades, but only recently have they become dramatically successful, as computational advances enabled the estimation of very large models, essentially aligning with the core idea of the universal approximation theorem. This has led finance and economics, as well as many other fields, to develop models with a very large number of parameters. The main shortcoming of neural networks for finance is that they are “black box” models. Once millions and sometimes hundred of millions of “factors” are run through often rather complicated and delicate gradient descent-like algorithms, it is difficult to see what features are economically important, how exactly the model produces results, and if an environment changes, whether the model still continues to perform. [^8] For a review of machine learning in asset pricing, see the book of Nagel (2021).
One alternative to neural nets is the Kolmogorov–Arnold representation theorem (Kolmogorov 1956; Arnold 1957) introduced in finance by Borri et al. (2024). [^9] Borri et al. (2024), arXiv, April 11, 2024. Kolmogorov–Arnold remarkably shows that any nonlinear function can be written exactly as a combination of a finite sum of univariate functions composed with a link function. This formulation is thus a representation, not an approximation. Importantly, the early machine learning literature showed how to use in theory this representation to construct a shallower and smaller network (Hecht-Nielsen 1987, Girosi & Poggio 1989, Kurkova 1991). This network can be thought of as a much lower dimensional neural network where the activation functions of each neuron are being trained and adapt to the data. While theoretically this result is remarkable, [^10] In fact, Kolmogorov and Arnold construction solved Hilbert’s thirteenth problem, one of the major open problems of mathematics of the twentieth century. it is not clear whether this approach would work on actual data because of the complexity that the theory requires for the representation of any function, and in particular, on finance data, which is both relatively small and has a very low signal-to-noise ratio. Borri et al. (2024) is the first paper that shows the practical implementation and effectiveness of the Kolmogorov–Arnold architecture on actual data. Specifically, they show that this methodology enables construction of a single non-linear factor (KA factor) that captures a significant part of the variation in expected returns. [^11] Schmidt-Hieber (2021) is the first paper in the literature that studies quantitative properties of Kolmogorov–Arnold representation for multiple-layer neural networks and thus introduces the Kolmogorov–Arnold architecture. An early paper by Horowitz & Mammen (2007) discusses essentially a similar structure of general nonparametric regressions with an unknown link function. Subsequent to Borri et al. (2024), Liu et al. (2024) (arXiv, April 30, 2024) popularized Kolmogorov Arnold Networks (KAN) to the machine learning community and showed how significantly shallower neural nets built on Kolmogorov–Arnold architecture may, in some cases, outperform and bring interpretability in several science applications. Toscano et al. (2025) argue that KAN rely on computation expensive basis functions with a performance that degrades with high parameter counts, which might limit the use in these science applications. The main advantage of this approach is that the KA-factor can be thought of as a simple representation of a much more complicated and larger neural network and, in principle, captures all such networks, without having to consider a variety of neural networks specifications and the optimization algorithms.
@@ 3.6 Fact 6: Few factors, higher orders—rather than machine learning: why less is more
Another alternative to neural networks follows the traditional finance toolbox—approximate and estimate the stochastic discount factor via a Taylor expansion. However, the Taylor approximation to the higher order generates a very large set of factors, due to the need to account for all of the interactions of the higher-order terms in the Taylor expansion. Borri et al. (2025) develop the properties of a forward-selection Fama–MacBeth (FS-FMB) method for selecting the higher-order factors and their interactions that are most important for the cross-section of returns. The method proceeds sequentially, adding one factor or interaction to the selected set at a time and, at each step, choosing the term that maximizes the Fama–MacBeth (FMB) second-step cross-sectional R-squared. The main advantage of this approach is that it is a fully “glass box” model both in terms of determining the important factors and in the process to select them that allows to map features of the data to the economic factors and their higher order interactions. [^12] Conceptually, these higher-order interaction terms capture conditional moments, as in systemic-risk measures such as CoVaR (Adrian & Brunnermeier, 2016) and marginal expected shortfall (Acharya et al., 2017). See Bell et al. (2024) for an alternative “glass box” machine learning structure.
Table 5, Panel A, reports second-step FMB results for the crypto-CAPM, or C-1, model and for the C-4 model of Liu et al. (2021). The test assets are the set of 50 portfolios formed using the smart beta strategies described in Fact 4 (see Tables 2 and 3), and the CARB strategy described in Fact 8 (see Table 6, Panel B). As with the classic CAPM, the C-1 model cannot account for the cross-section of average returns. By contrast, the C-4 model explains 47.3% of the cross-sectional variation, and satisfies the restriction of a zero intercept.
Table 5, Panel B, reports the result of a specification which augments the C-4 model with the single KA-factor based on the Kolmogorov-Arnold architecture of Borri et al. (2024). One can think of this factor as a stand-in that captures a variety of different neural nets. The cross-sectional R-squared increases by approximately 25 pp, and the intercept is not statistically different from zero at standard levels.
@@
Table 5, Panel C, reports the results from the estimation of the FS-FMB (Borri et al. 2025). At each step, we augment the C-4 by one additional higher-order factor ($h_{j}$) which maximizes the adjusted cross-sectional R-squared. Higher-order factors include the squared and cubed of the factors in the C-4 model, as well their pair-wise interactions up to order three. The procedure selects three higher-order factors. The first is CMOM cubed, resulting in an increase of the cross-sectional R-squared of approximately 15 pp. The importance of this factor is that it relates to the frequent and large jumps which are the distinct feature of cryptocurrency of Fact 5. The second and third selected factors are CMKT squared and the interaction between CMKT and CSIZE. Borri et al. (2025) show that the equity counterparts of these factors are important for the cross-section of equity returns. The cross-sectional R-squared of the FS-FMB procedure is 70%, a value very similar to that of the model including the KA-factor. In all models, the estimate of the intercept is not statistically different from zero.
@@
These results show that the FS–FMB procedure recovers essentially the same nonlinear structure as the KA single factor. Moreover, since KA can represent any neural net, FS–FMB provides a simple, transparent, “open box” alternative to black-box models that is directly connected to the economic structure of the model and to observables characteristics. In short, rather than estimating a neural network (or studying a particular variant of the multitude of neural networks), one can estimate the FS–FMB specification and capture the relevant nonlinearities in a parsimonious and interpretable way without relying on millions of parameters.
@@
This table reports the results from the second step of the FMB method. Panel A refers to baseline models: the crypto-CAPM (C-1) and the C-4 model of Liu et al. (2022). The table reports the adjusted cross-sectional R-squared, the estimate for the intercept ($\alpha$) and associated t-statistic. The t-statistics is based on Newey-West corrected standard errors. Panel B refers to the C-4 model augmented by the single KA-factor, estimated with the KA procedure of Borri et al. (2024). Panel C refers to the FS-FMB procedure. At each step, we augment the C-4 model by one additional higher-order factor ($h_{j}$) which maximizes the adjusted cross-sectional R-squared. The method stops when the improvement in the R-squared in a given step is smaller than 1 pp.
3.7 Fact 7: In crypto, the (block)chain drives the gain
Fact 6: Blockchain economics and information significantly affects cryptocurrency prices, on-chain user adoption accounts for about 8% of the variation in cryptocurrency returns.
Network externalities are a fundamental concept in economics because they shape markets and influence competition, regulation and technology adoption. Classic work has analyzed how networks form and evolve (Jackson & Wolinsky, 1996), how pricing and competition operate on networked platforms (Rochet & Tirole, 2003), and how shocks propagate through financial networks (Allen & Gale, 2000). A central question common in this work is information diffusion and learning in a network (Golub & Jackson 2010, Acemoglu et al. 2011).

### v3

@@ 3.2 Stylized fact 2: How to be “smart” in crypto—crypto-size, crypto-momentum and crypto-value
Stylized fact 2: A small number of smart beta strategies—based on crypto-size, crypto-momentum, and crypto-value—span the cross-section of cryptocurrency returns.
@@
A central question in asset pricing is whether systematic variation in expected returns can be explained by a parsimonious set of factors. In equities, substantial academic endeavor—most prominently by Eugene Fama and Kenneth French (e.g., Fama & French 1992)—established this approach, and factor models play a core role in modern finance practice and theory from risk management of large banks to niche hedge fund strategies. For cryptocurrencies, Liu, Tsyvinski & Wu (2022) establish that a three factor model consisting of the aggregate cryptocurrency market, size, and momentum (C-3 factor model), captures a large fraction of the cross-section of expected returns. Liu, Tsyvinski & Wu (2021) show that adding a value factor (C-4 factor model) further strengthens the explanatory power. We revisit this evidence using our updated sample and show that a small number of predictors drive the cross-section of cryptocurrency returns over the longer sample and post-2020.
@@
We construct smart beta strategies by sorting coins into quintile portfolios based on three established significant predictors: market capitalization (CSIZE), past two-week returns (CMOM), and the price-to-new-address ratio (CVALUE). [^9] Studies based on smaller samples of coins have also proposed additional predictors of cryptocurrency returns in the cross-section, including liquidity (e.g., Bianchi, Babiak & Dickerson 2022), market segmentation (e.g., Borri & Shakhnov 2022), and higher-order moments (e.g., Borri et al. 2022b; Borri & Santucci de Magistris 2022). The return to each predictor is measured as the long-short spread between the highest and lowest quintiles (Portfolio 5 minus Portfolio 1).
@@
Table 2, Panels A and B, report average excess returns to the quintile portfolios and the corresponding long-short strategies for the full sample and the post-2020 subsample. All three predictors generate statistically significant long-short returns in the full sample. The average long-short returns based on size and value are significantly negative and the average long-short return based on momentum is significantly positive, consistent with Liu, Tsyvinski & Wu (2022). Cryptocurrency value is the weekly quintile portfolio of each bracket based on the price-to-new address ratio. High bracket is the low-value coins and low bracket is the high-value coins. The sorting is done in week $t$ and the returns are the value-weighted average in the following week.
@@
The C-4 structure of the crypto-market thus parallels the equity market, where a small number of equity-market factors—most notably size, value, momentum, and profitability—dominate the cross-section of expected returns. For example, Asness, Moskowitz & Pedersen (2013) present evidence of value and momentum in equities, bonds, currencies, and commodities. The C-4 factor model of Liu, Tsyvinski & Wu (2022) and Liu, Tsyvinski & Wu (2021) is the cryptocurrency counterpart of the classic 4-factor model of Fama & French (1992) and Carhart (1997) for equities. It is important to note that the predictors in cryptocurrency markets are only partly analogous to those in equity markets. CSIZE and CMOM closely resemble their equity counterparts, although crypto momentum is measured over a much shorter horizon (previous two weeks, versus the standard 12-to-2 months in equities). In contrast, there is no direct analogue to book value (Shiller 1981) in cryptocurrencies, so we follow Liu, Tsyvinski & Wu (2021) and proxy CVALUE with the price-to-new-address ratio. [^10] Liu, Tsyvinski & Wu (2021) show that the number of new addresses captures the network effects of cryptocurrencies and serves as a measure of their fundamental value.
Establishing the small factor structure of the cryptocurrency returns, in our view, is not only important from purely the empirical asset pricing point of view but also since it provides a set of stylized low dimensional features that capture the main insights of several important theoretical models of cryptocurrency (Cong, Li & Wang 2021b; Sockin & Xiong 2023a, b). More broadly, this stylized low-dimensional structure is consistent with the broader multifactor view of asset pricing developed for equities, but the relevant priced risks in cryptocurrency include distinct, crypto-specific components—including network/adoption-related fundamentals.
@@
A related important point is that in addition to determining which betas are smart, it is equally as important to establish which betas from the factor zoo are not “smart”. Specifically, Liu, Tsyvinski & Wu (2022) and Liu, Tsyvinski & Wu (2021) show that—besides crypto-specific size, momentum, and value—most of the smart beta strategies from the canonical factor zoo in the equity market (Feng, Giglio & Xiu 2020) either do not generate significant hedged long-short strategy returns or are subsumed by the C-4 factor model. In Table 3, we study long-short strategies based on these signals post-2020. We find that these strategies continue not to generate significant average returns in the recent sample. [^11] In a recent study, Fieberg et al. (2024) examine 20,736 research designs for 43 sorting variables in the cryptocurrency market. They find that the most prominent cryptocurrency factors, such as size and momentum, remain consistently robust across various specifications, but the others do not. These results show that the cross-section of cryptocurrency returns can be summarized by a small number of factors.
@@
Notes: This table reports the mean quintile portfolio returns and long-short portfolio return for crypto smart beta strategies based on insignificant predictors. The predictors are past 12-week and 24-week return (MOM12 and MOM24), past 24-week return standard deviation (VOL), past 24-week dollar trading volume (VOLUME), coin market beta (CMKT BETA) and stock market beta (MKT BETA) estimated using past 24-week returns. The long-short portfolio is long portfolio 5 and short portfolio 1 for all predictors. The mean returns are the time-series averages of weekly value-weighted portfolio excess returns. The post-2020 sample is the sample since Jan 1, 2020. $t$-statistics in parentheses are Newey-West adjusted. *, **, and *** denote significance at the 10%, 5% and 1% levels.
3.3 Stylized fact 3: Mind the Jumps—few factors, higher orders: why less is more
Stylized fact 3: Cryptocurrency returns feature frequent jumps; in the cross-section, a parsimonious model augmented with a small number of higher-order terms captures most of the predictive content of black-box machine learning.
@@ 3.3 Stylized fact 3: Mind the Jumps—few factors, higher orders: why less is more
Leibniz wrote that natura non facit saltus—nature makes no jumps. In economics and finance, this often means that most changes occur gradually, with only occasional discontinuities. Indeed, jumps are rare in many asset classes, and this key insight leads Black & Scholes (1973) to model prices as Brownian motions. Cryptocurrencies are different: large, sudden price moves are common. Borri & Santucci de Magistris (2022) show that jumps account for a sizable share of Bitcoin’s return variation and that higher-order moments help explain its risk premium. [^12] See early evidence from Scaillet, Treccani & Trevisan (2020) similarly documenting very frequent jumps in Bitcoin.
Table 4 reports the frequency and magnitude of statistically significant jumps in Bitcoin and Ethereum—relative to U.S. equities—using realized-variance and bipower-variation measures constructed from 5-minute data and the BNS test. Jumps occur far more frequently in cryptocurrencies than in the Dow Jones stocks, although conditional on a jump the share of variation attributable to the jump is broadly similar across assets. In short, contrary to Leibniz’s dictum, in crypto markets saltus sunt regula (jumps are the rule), not the exception.
@@
This table reports the probability of observing a statistically significant jump (P(J)), the number of such events (N(J)), and quantiles of the ratio of jump variation to realized variance (J/RV), conditional on a significant jump, for Bitcoin, Ethereum, and U.S. stocks. Jumps are identified as the difference between the daily realized variance of Barndorff-Nielsen & Shephard (2002) and the bi-power variation of Barndorff-Nielsen & Shephard (2006). The two realized measures are constructed from tick-level data aggregated at the 5-minute frequency. Significant jumps are identified using the BNS test with a 0.1% significance level. For U.S. equities, values are cross-sectional averages across the constituents of the Dow Jones Industrial Average. Data are from January 1, 2015 to March 31, 2025. The pre-2020 sample is the sample ending Jan 1, 2020. The post-2020 sample is the sample since Jan 1, 2020.
This prevalence of jumps, and the corresponding role of higher-order moments, poses challenges for risk management and portfolio choice (e.g., Borri 2019) and has a direct implication for asset pricing: nonlinearities matter. Motivated by this, a growing literature applies machine-learning methods—especially neural networks—to the cross-section of returns (e.g., Gu, Kelly & Xiu 2020). While neural networks have a universal approximation property (Cybenko 1989; Hornik, Stinchcombe & White 1989), their main drawback in finance is that they are “black box” models: they can fit complex patterns but make it difficult to isolate which economic features drive predicted returns and whether those relationships are stable when the environment changes. [^13] For a review of machine learning in asset pricing, see Nagel (2021).
One alternative to neural nets is the Kolmogorov-Arnold (KA) representation theorem (Kolmogorov 1956; Arnold 1957), introduced in finance by Borri et al. (2024). [^14] Borri et al. (2024), arXiv, April 11, 2024. The theorem shows that a broad class of nonlinear functions can be exactly represented as finite sums of univariate functions composed with a link function—a structure that motivates shallow, low-dimensional architectures (see, e.g., Hecht-Nielsen 1987; Girosi & Poggio 1989; Kurkova 1991). In practice, Borri et al. (2024) implement this idea and construct a single nonlinear factor (the KA factor) that captures a sizable share of the cross-sectional variation in expected returns. [^15] Related work includes Schmidt-Hieber (2021) and Horowitz & Mammen (2007). Subsequent to Borri et al. (2024), Liu et al. (2024) popularize KAN in the machine-learning community, while Toscano, Wang & Karniadakis (2025) discuss computational and scaling limitations.
@@
Another alternative to neural networks follows the traditional finance toolbox—approximate and estimate the stochastic discount factor via a Taylor expansion. However, the Taylor approximation to the higher order generates a very large set of factors, including all of the interactions of the higher-order terms in the Taylor expansion. Borri et al. (2025) develop the properties of a forward-selection Fama-MacBeth (FS-FMB) method for selecting the higher-order factors and their interactions that are most important for the cross-section of returns by sequentially adding one factor or interaction to the selected set at a time and, at each step, choosing the term that maximizes the Fama–MacBeth (FMB) second-step cross-sectional R-squared. This approach is a fully “glass box” model in terms of determining the important factors. [^16] Conceptually, these higher-order interaction terms capture conditional moments, as in systemic-risk measures such as CoVaR (Adrian & Brunnermeier, 2016) and marginal expected shortfall (Acharya et al., 2017). See Bell et al. (2024) for an alternative “glass box” machine learning structure.
@@
Table 5, Panel A, reports second-step FMB results for the crypto-CAPM, or C-1, model and for the C-4 model of Liu, Tsyvinski & Wu (2021). The test assets are the set of 50 portfolios formed using the smart beta strategies described in stylized fact 2 (see Tables 2 and 3), and the CARB strategy described in stylized fact 5 (see Table 6, Panel B). As with the classic CAPM, the C-1 model cannot account for the cross-section of average returns. By contrast, the C-4 model explains 47.3% of the cross-sectional variation, and satisfies the restriction of a zero intercept. Table 5, Panel B, reports the result of a specification which augments the C-4 model with the single KA-factor based on the Kolmogorov-Arnold architecture of Borri et al. (2024). One can think of this factor as a stand-in that captures a variety of different neural nets. The cross-sectional R-squared increases by approximately 25 pp, and the intercept is not statistically different from zero at standard levels. Table 5, Panel C, reports the results from the estimation of the FS-FMB (Borri et al. 2025). At each step, we augment the C-4 by one additional higher-order factor ($h_{j}$) which maximizes the adjusted cross-sectional R-squared. Higher-order factors include the squared and cubed of the factors in the C-4 model, as well their pair-wise interactions up to order three. The procedure selects three higher-order factors. The cross-sectional R-squared of the FS-FMB procedure is 70%, a value very similar to that of the model including the KA-factor. In all models, the estimate of the intercept is not statistically different from zero.
@@
These results show that the FS-FMB procedure recovers essentially the same nonlinear structure as the KA single factor. Moreover, since KA can represent any neural net, FS–FMB provides a simple, transparent, “open box” alternative to black-box models that is directly connected to the economic structure of the model and to observables characteristics. In short, rather than estimating a neural network (or studying a particular variant of the multitude of neural networks), one can estimate the FS–FMB specification and capture the relevant nonlinearities in a parsimonious and interpretable way without relying on millions of parameters.
@@
This table reports the results from the second step of the FMB method. Panel A refers to baseline models: the crypto-CAPM (C-1) and the C-4 model of Liu, Tsyvinski & Wu (2022). The table reports the adjusted cross-sectional R-squared, the estimate for the intercept ($\alpha$) and associated t-statistic. The t-statistics are based on Newey-West corrected standard errors. Panel B refers to the C-4 model augmented by the single KA-factor, estimated with the KA procedure of Borri et al. (2024). Panel C refers to the FS-FMB procedure. At each step, we augment the C-4 model by one additional higher-order factor ($h_{j}$) which maximizes the adjusted cross-sectional R-squared. The method stops when the improvement in the R-squared in a given step is smaller than 1 pp.
3.4 Stylized fact 4: In crypto, the (block)chain drives the gain
Stylized fact 4: Blockchain economics and information significantly affect cryptocurrency prices, on-chain user adoption accounts for about 8% of the variation in cryptocurrency returns.
Network externalities are a fundamental concept in economics because they shape markets and influence competition, regulation and technology adoption. Classic work has analyzed how networks form and evolve (Jackson & Wolinsky, 1996), how pricing and competition operate on networked platforms (Rochet & Tirole, 2003), and how shocks propagate through financial networks (Allen & Gale, 2000). A central question common in this work is information diffusion and learning in a network (Golub & Jackson 2010; Acemoglu et al. 2011).